# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - Blockchain and Applications

### v2

@@ 2.3.1 Fundamentals of Block Cipher Encryption
Block ciphers are encryption methods that operate on fixed-size blocks of data[59]. In block ciphers, the input data is divided into blocks of a specified length (typically 64 or 128 bits), and each block is encrypted separately using a symmetric key. This is in contrast to stream ciphers, which encrypt data bit by bit or byte by byte in a continuous stream.
@@ 2.3.2 Block Cipher Modes: ECB, CBC, CFB, OFB, CTR
In ECB mode[60], each block of plaintext is encrypted independently of other blocks. While this mode is simple and fast, it has a significant weakness: identical plaintext blocks produce identical ciphertext blocks. This can reveal patterns in the plaintext, making ECB unsuitable for encrypting large datasets or images.
@@
CBC mode addresses the weaknesses of ECB by chaining the blocks together[61]. In this mode, each plaintext block is XORed with the previous ciphertext block before encryption. To ensure the first block is encrypted securely, an initialization vector (IV) is used. CBC is more secure than ECB but requires sequential encryption, making parallel processing difficult.
@@
CFB operates similarly to CBC, but it can process smaller units than the block size, such as bits or bytes, making it more flexible for streaming data[62]. In CFB, the previous ciphertext block is encrypted, and the result is XORed with the plaintext to produce the current ciphertext block. It also requires an IV but allows error propagation.
@@
OFB mode is similar to CFB but avoids error propagation[63]. It generates a keystream by repeatedly encrypting the IV, which is XORed with the plaintext to produce the ciphertext. Since the plaintext never directly influences the ciphertext, a transmission error in one block does not affect subsequent blocks.
@@
In CTR mode, instead of chaining blocks, a counter is used, which is incremented for each block[64]. The counter is encrypted, and the result is XORed with the plaintext to produce the ciphertext. CTR mode supports parallel processing of blocks and is efficient for applications where fast encryption is needed.
@@ 2.4.1 Caesar Cipher
The Caesar cipher[65] is one of the simplest and oldest encryption algorithms. It is a substitution cipher named after Julius Caesar, who reportedly used it to protect military communications. The Caesar cipher works by shifting each letter in the plaintext by a fixed number of positions in the alphabet. For example, with a shift of 3, A becomes D, B becomes E, and so on. Once the end of the alphabet is reached, the cipher wraps around to the beginning.
@@ 2.4.2 Vigenère Cipher
The Vigenère cipher is a polyalphabetic substitution cipher[66], which means it uses multiple substitution alphabets to encode the plaintext. In contrast to a simple substitution cipher, where each letter in the plaintext is always replaced by the same letter in the ciphertext, the Vigenère cipher varies the substitution depending on the position of the letter and the key.
@@
The Kasiski examination[67] is a method used to break the Vigenère cipher by identifying repeated sequences in the ciphertext. Since the cipher relies on a repeating keyword, patterns can emerge when the key is repeated, especially if parts of the plaintext contain repeated words or phrases.
@@ 2.4.3 Data Encryption Standard (DES)
The Data Encryption Standard (DES) is a symmetric-key block cipher that was widely used for data encryption in the late 20th century[68]. DES operates on blocks of 64 bits of data and uses a 56-bit key. The core of DES is a Feistel network[69], which involves splitting the data into two halves and applying a series of encryption rounds.
@@
Furthermore, DES is vulnerable to differential cryptanalysis and linear cryptanalysis, which exploit specific patterns in the encryption process. Due to these weaknesses, DES was replaced by more secure algorithms, such as Triple DES (3DES)[70] and Advanced Encryption Standard (AES)[71].
@@ 2.4.4 Triple DES (3DES)
Triple DES (3DES)[72] enhances the security of the original Data Encryption Standard (DES) by applying the DES algorithm three times with three different keys. The basic operation process is as follows:
@@ 2.4.5 Advanced Encryption Standard (AES)
The Advanced Encryption Standard (AES)[71] is a symmetric block cipher that was developed as a replacement for DES and 3DES. AES was selected as the encryption standard by the U.S. government in 2001[73]. Its design is based on a structure known as the Substitution-Permutation Network (SPN)[74], which is resistant to several forms of cryptanalysis, including linear and differential cryptanalysis.
@@
  • Substitution: AES uses a non-linear substitution step (called the S-box)[75] that replaces each byte with another byte according to a fixed lookup table. This helps introduce confusion into the data, making it harder for attackers to find relationships between the plaintext and ciphertext.
  • Permutation: AES also includes a permutation step (known as the ShiftRows and MixColumns steps)[76], which re-arranges the data to further complicate the relationships between the plaintext and ciphertext. This adds diffusion, spreading the influence of each byte of the plaintext across the entire ciphertext.
@@ 3.1 Basic Concepts of Asymmetric Encryption
Asymmetric encryption, also known as public-key encryption, is a cryptographic method that uses a pair of keys to encrypt and decrypt data[38]. This approach eliminates the need for both parties to share a secret key ahead of time, making communication more secure.
@@ 3.2 RSA Algorithm
The RSA algorithm[77] is one of the most widely used asymmetric encryption algorithms. It is based on the difficulty of factoring large numbers, which ensures the security of the cryptosystem.
@@ 3.3.1 Mathematical Foundation of ECC
Elliptic Curve Cryptography (ECC)[78] is based on the algebraic structure of elliptic curves over finite fields. An elliptic curve is represented by an equation of the form:
@@
One of the most important properties of elliptic curves is the difficulty of the Elliptic Curve Discrete Logarithm Problem (ECDLP)[79]. Given two points $P$ and $Q$ on an elliptic curve, where $Q=kP$ (i.e., $Q$ is the result of adding $P$ to itself $k$ times), finding $k$ is computationally hard. This problem provides the security foundation for ECC, similar to how the difficulty of factoring large numbers provides the security for RSA.
@@ 3.3.4 Common Applications of ECC
  • Mobile Device Encryption: ECC is used in mobile devices to secure communications and data[80]. Due to the limited computational power and battery life of mobile devices, ECC’s efficiency makes it a preferred choice for securing messaging apps, email, and VPNs.
  • Cryptocurrency (e.g., Bitcoin): ECC plays a crucial role in securing cryptocurrencies like Bitcoin[81]. The Bitcoin protocol uses ECC to generate cryptographic keys and sign transactions, ensuring that only the owner of the private key can authorize spending.
  • SSL/TLS for Secure Web Communication: ECC is used in SSL/TLS protocols to secure web communications[82]. When you visit websites with HTTPS, ECC may be part of the underlying cryptography that secures your connection.
@@ 3.4.1 Working Mechanism of Diffie-Hellman
The Diffie-Hellman key exchange[83] is a fundamental cryptographic protocol that allows two parties to securely generate a shared secret over an insecure communication channel. This shared secret can then be used for symmetric encryption of further communication. The beauty of this protocol is that it does not require the two parties to have previously exchanged any secret keys. Instead, they exchange some public values, from which the shared secret is derived.
@@ 3.4.2 Security Analysis of the Key Exchange Process
The security of the Diffie-Hellman key exchange protocol relies heavily on the difficulty of solving the discrete logarithm problem[84]. Even if an eavesdropper, often referred to as Eve, can observe the public values exchanged (the prime $p$, the base $g$, and the public keys $A$ and $B$), calculating the shared secret requires solving $g^{a}\mod p$ or $g^{b}\mod p$, which is practically impossible for large $p$.
@@
Vulnerability to Man-in-the-Middle (MITM) Attacks: However, Diffie-Hellman is vulnerable to a man-in-the-middle attack if the public keys are not authenticated[85]. In this attack, an adversary, often called ”Mallory,” intercepts the public key exchange between Party A and Party B. Mallory can then generate her own private key and public key, replacing Party A’s public key with her own in the message to Party B and vice versa. This allows Mallory to establish two separate shared secrets with each party, enabling her to decrypt and re-encrypt all communications between them without their knowledge.
To prevent man-in-the-middle attacks, it is essential to authenticate the public keys, typically through the use of certificates or digital signatures. Modern implementations of Diffie-Hellman, such as those in Transport Layer Security (TLS)[86], often use authenticated variants to address this issue.
@@ 3.4.3 Modern Applications of Diffie-Hellman
  • Virtual Private Networks (VPNs): VPNs often use Diffie-Hellman to establish a secure connection between the client and the server[87]. The shared secret derived from Diffie-Hellman is used to encrypt the data transmitted between the two endpoints.
  • TLS/SSL (Transport Layer Security): Secure web communications, such as HTTPS, use a variant of Diffie-Hellman, known as Elliptic Curve Diffie-Hellman (ECDH)[88], to negotiate encryption keys during the handshake process. This ensures that even if the communication channel is intercepted, the eavesdropper cannot decrypt the data.
  • Secure Messaging: Protocols like the Signal Protocol use Diffie-Hellman to provide forward secrecy in encrypted messaging applications[89]. This ensures that even if a user’s encryption keys are compromised in the future, previous communications remain secure.
@@ 3.5.1 Comparison of Encryption Efficiency
Symmetric encryption and asymmetric encryption differ significantly in terms of their efficiency. Symmetric encryption algorithms, such as AES (Advanced Encryption Standard), are much faster than asymmetric algorithms like RSA (Rivest-Shamir-Adleman)[40]. This is because symmetric encryption uses the same key for both encryption and decryption, allowing for efficient operations, typically involving block ciphers or stream ciphers.
@@ 3.5.3 Use Cases of Asymmetric Encryption in Blockchain
  • Digital Signatures: In blockchain systems like Bitcoin and Ethereum, transactions are signed using the sender’s private key[90]. This ensures that only the holder of the private key can initiate a transaction. The public key can be used by others to verify the authenticity of the transaction.
@@
  • Securing Transactions: Cryptographic techniques based on asymmetric encryption, such as elliptic curve cryptography (ECC), are used to secure transactions in modern blockchain platforms[91]. These techniques ensure that transactions are tamper-proof and can be validated by all participants in the network.
@@ 4.1 Basic Definition and Properties of Hash Functions
Hash functions play a fundamental role in many cryptographic systems, particularly in blockchain technology. At their core, a hash function is a mathematical algorithm that takes an input (or ”message”) and returns a fixed-size string of bytes, typically in the form of a hash value[92]. This output is unique to the specific input, but the same input will always result in the same hash value. For example, in blockchain systems, hash functions are used to maintain data integrity, verify transactions, and protect the system from manipulation. There are several essential properties of hash functions, two of which are particularly important: one-wayness and collision resistance.
@@ 4.1.1 One-Wayness and Irreversibility
One of the most critical properties of a cryptographic hash function is one-wayness[93]. This means that, given the hash value, it is computationally infeasible to reverse-engineer the input that produced it. Essentially, hash functions are designed to be irreversible. Even if someone has the hash value, they cannot retrieve the original input data easily.
@@
This irreversibility is what protects data in blockchain systems[94]. For example, even if someone obtains the hash of a block or transaction, they cannot use it to figure out the details of the original transaction.
@@ 4.1.2 Collision Resistance and Security
Another critical property of cryptographic hash functions is collision resistance[95]. A collision occurs when two different inputs produce the same hash output. A secure hash function is designed to make it computationally infeasible to find any two distinct inputs that result in the same hash value.
@@ 4.2 Simple Hash Algorithms
While cryptographic hash functions are designed with complex mathematics to ensure security, there are simpler forms of hash functions that are used in various applications, especially when high security is not the main concern. These simpler algorithms are often used for tasks[96] such as quick data indexing, load balancing, or data distribution across different storage locations.
@@ 4.2.1 Modulo Operation and Hashing
One of the most straightforward ways to generate a hash value is by using the modulo operation[97]. In this approach, the hash value is obtained by dividing the input value by a certain number (called the modulus) and taking the remainder.

### v3

@@ 2.3.1 Fundamentals of Block Cipher Encryption
Block ciphers are encryption methods that operate on fixed-size blocks of data[Singh2014BlockCI]. In block ciphers, the input data is divided into blocks of a specified length (typically 64 or 128 bits), and each block is encrypted separately using a symmetric key. This is in contrast to stream ciphers, which encrypt data bit by bit or byte by byte in a continuous stream.
@@ 2.3.2 Block Cipher Modes: ECB, CBC, CFB, OFB, CTR
In ECB mode[Rachman2010PERBANDINGANMC], each block of plaintext is encrypted independently of other blocks. While this mode is simple and fast, it has a significant weakness: identical plaintext blocks produce identical ciphertext blocks. This can reveal patterns in the plaintext, making ECB unsuitable for encrypting large datasets or images.
@@
CBC mode addresses the weaknesses of ECB by chaining the blocks together[Vaidehi2014DesignAA]. In this mode, each plaintext block is XORed with the previous ciphertext block before encryption. To ensure the first block is encrypted securely, an initialization vector (IV) is used. CBC is more secure than ECB but requires sequential encryption, making parallel processing difficult.
@@
CFB operates similarly to CBC, but it can process smaller units than the block size, such as bits or bytes, making it more flexible for streaming data[Gope2015IntegrityAwarePC]. In CFB, the previous ciphertext block is encrypted, and the result is XORed with the plaintext to produce the current ciphertext block. It also requires an IV but allows error propagation.
@@
OFB mode is similar to CFB but avoids error propagation[Sung2001ConcreteSA]. It generates a keystream by repeatedly encrypting the IV, which is XORed with the plaintext to produce the ciphertext. Since the plaintext never directly influences the ciphertext, a transmission error in one block does not affect subsequent blocks.
@@
In CTR mode, instead of chaining blocks, a counter is used, which is incremented for each block[Rogaway2012CommentsTN]. The counter is encrypted, and the result is XORed with the plaintext to produce the ciphertext. CTR mode supports parallel processing of blocks and is efficient for applications where fast encryption is needed.
@@ 2.4.1 Caesar Cipher
The Caesar cipher[Allen2017TheCC] is one of the simplest and oldest encryption algorithms. It is a substitution cipher named after Julius Caesar, who reportedly used it to protect military communications. The Caesar cipher works by shifting each letter in the plaintext by a fixed number of positions in the alphabet. For example, with a shift of 3, A becomes D, B becomes E, and so on. Once the end of the alphabet is reached, the cipher wraps around to the beginning.
@@ 2.4.2 Vigenère Cipher
The Vigenère cipher is a polyalphabetic substitution cipher[RubinsteinSalzedo2018TheVC], which means it uses multiple substitution alphabets to encode the plaintext. In contrast to a simple substitution cipher, where each letter in the plaintext is always replaced by the same letter in the ciphertext, the Vigenère cipher varies the substitution depending on the position of the letter and the key.
@@
The Kasiski examination[Hananto2019AnalyzingTK] is a method used to break the Vigenère cipher by identifying repeated sequences in the ciphertext. Since the cipher relies on a repeating keyword, patterns can emerge when the key is repeated, especially if parts of the plaintext contain repeated words or phrases.
@@ 2.4.3 Data Encryption Standard (DES)
The Data Encryption Standard (DES) is a symmetric-key block cipher that was widely used for data encryption in the late 20th century[Landau2000StandingTT]. DES operates on blocks of 64 bits of data and uses a 56-bit key. The core of DES is a Feistel network[Kumari2023AnRI], which involves splitting the data into two halves and applying a series of encryption rounds.
@@
Furthermore, DES is vulnerable to differential cryptanalysis and linear cryptanalysis, which exploit specific patterns in the encryption process. Due to these weaknesses, DES was replaced by more secure algorithms, such as Triple DES (3DES)[Venigalla2012IMPLEMENTATIONOT] and Advanced Encryption Standard (AES)[Selent2010ADVANCEDES].
@@ 2.4.4 Triple DES (3DES)
Triple DES (3DES)[Chen2021AccountingDE] enhances the security of the original Data Encryption Standard (DES) by applying the DES algorithm three times with three different keys. The basic operation process is as follows:
@@ 2.4.5 Advanced Encryption Standard (AES)
The Advanced Encryption Standard (AES)[Selent2010ADVANCEDES] is a symmetric block cipher that was developed as a replacement for DES and 3DES. AES was selected as the encryption standard by the U.S. government in 2001[Rao2017ASO]. Its design is based on a structure known as the Substitution-Permutation Network (SPN)[Li2011AND], which is resistant to several forms of cryptanalysis, including linear and differential cryptanalysis.
@@
  • Substitution: AES uses a non-linear substitution step (called the S-box)[Scripcariu2013OnTS] that replaces each byte with another byte according to a fixed lookup table. This helps introduce confusion into the data, making it harder for attackers to find relationships between the plaintext and ciphertext.
  • Permutation: AES also includes a permutation step (known as the ShiftRows and MixColumns steps)[Mahrousa2020ANM], which re-arranges the data to further complicate the relationships between the plaintext and ciphertext. This adds diffusion, spreading the influence of each byte of the plaintext across the entire ciphertext.
@@ 3.1 Basic Concepts of Asymmetric Encryption
Asymmetric encryption, also known as public-key encryption, is a cryptographic method that uses a pair of keys to encrypt and decrypt data[Simmons1979SymmetricAA]. This approach eliminates the need for both parties to share a secret key ahead of time, making communication more secure.
@@ 3.2 RSA Algorithm
The RSA algorithm[Zhong2022AnOO] is one of the most widely used asymmetric encryption algorithms. It is based on the difficulty of factoring large numbers, which ensures the security of the cryptosystem.
@@ 3.3.1 Mathematical Foundation of ECC
Elliptic Curve Cryptography (ECC)[Koppl2021ApplicationOC] is based on the algebraic structure of elliptic curves over finite fields. An elliptic curve is represented by an equation of the form:
@@
One of the most important properties of elliptic curves is the difficulty of the Elliptic Curve Discrete Logarithm Problem (ECDLP)[Sadkhan2021DevelopmentOS]. Given two points $P$ and $Q$ on an elliptic curve, where $Q=kP$ (i.e., $Q$ is the result of adding $P$ to itself $k$ times), finding $k$ is computationally hard. This problem provides the security foundation for ECC, similar to how the difficulty of factoring large numbers provides the security for RSA.
@@ 3.3.4 Common Applications of ECC
  • Mobile Device Encryption: ECC is used in mobile devices to secure communications and data[Dar2021LightweightAS]. Due to the limited computational power and battery life of mobile devices, ECC’s efficiency makes it a preferred choice for securing messaging apps, email, and VPNs.
  • Cryptocurrency (e.g., Bitcoin): ECC plays a crucial role in securing cryptocurrencies like Bitcoin[Mehta2015ASO]. The Bitcoin protocol uses ECC to generate cryptographic keys and sign transactions, ensuring that only the owner of the private key can authorize spending.
  • SSL/TLS for Secure Web Communication: ECC is used in SSL/TLS protocols to secure web communications[Sebastian2015AdvantageOU]. When you visit websites with HTTPS, ECC may be part of the underlying cryptography that secures your connection.
@@ 3.4.1 Working Mechanism of Diffie-Hellman
The Diffie-Hellman key exchange[Kumar2020DiffieHS] is a fundamental cryptographic protocol that allows two parties to securely generate a shared secret over an insecure communication channel. This shared secret can then be used for symmetric encryption of further communication. The beauty of this protocol is that it does not require the two parties to have previously exchanged any secret keys. Instead, they exchange some public values, from which the shared secret is derived.
@@ 3.4.2 Security Analysis of the Key Exchange Process
The security of the Diffie-Hellman key exchange protocol relies heavily on the difficulty of solving the discrete logarithm problem[GanjewarDieHK]. Even if an eavesdropper, often referred to as Eve, can observe the public values exchanged (the prime $p$, the base $g$, and the public keys $A$ and $B$), calculating the shared secret requires solving $g^{a}\mod p$ or $g^{b}\mod p$, which is practically impossible for large $p$.
@@
Vulnerability to Man-in-the-Middle (MITM) Attacks: However, Diffie-Hellman is vulnerable to a man-in-the-middle attack if the public keys are not authenticated[Naher2018AuthenticationOD]. In this attack, an adversary, often called ”Mallory,” intercepts the public key exchange between Party A and Party B. Mallory can then generate her own private key and public key, replacing Party A’s public key with her own in the message to Party B and vice versa. This allows Mallory to establish two separate shared secrets with each party, enabling her to decrypt and re-encrypt all communications between them without their knowledge.
To prevent man-in-the-middle attacks, it is essential to authenticate the public keys, typically through the use of certificates or digital signatures. Modern implementations of Diffie-Hellman, such as those in Transport Layer Security (TLS)[Eronen2005PreSharedKC], often use authenticated variants to address this issue.
@@ 3.4.3 Modern Applications of Diffie-Hellman
  • Virtual Private Networks (VPNs): VPNs often use Diffie-Hellman to establish a secure connection between the client and the server[Carts2001ARO]. The shared secret derived from Diffie-Hellman is used to encrypt the data transmitted between the two endpoints.
  • TLS/SSL (Transport Layer Security): Secure web communications, such as HTTPS, use a variant of Diffie-Hellman, known as Elliptic Curve Diffie-Hellman (ECDH)[BlakeWilson2006EllipticCC], to negotiate encryption keys during the handshake process. This ensures that even if the communication channel is intercepted, the eavesdropper cannot decrypt the data.
  • Secure Messaging: Protocols like the Signal Protocol use Diffie-Hellman to provide forward secrecy in encrypted messaging applications[CohnGordon2018OnEE]. This ensures that even if a user’s encryption keys are compromised in the future, previous communications remain secure.
@@ 3.5.1 Comparison of Encryption Efficiency
Symmetric encryption and asymmetric encryption differ significantly in terms of their efficiency. Symmetric encryption algorithms, such as AES (Advanced Encryption Standard), are much faster than asymmetric algorithms like RSA (Rivest-Shamir-Adleman)[Raigoza2016EvaluatingPO]. This is because symmetric encryption uses the same key for both encryption and decryption, allowing for efficient operations, typically involving block ciphers or stream ciphers.
@@ 3.5.3 Use Cases of Asymmetric Encryption in Blockchain
  • Digital Signatures: In blockchain systems like Bitcoin and Ethereum, transactions are signed using the sender’s private key[Mayer2016ECDSASI]. This ensures that only the holder of the private key can initiate a transaction. The public key can be used by others to verify the authenticity of the transaction.
@@
  • Securing Transactions: Cryptographic techniques based on asymmetric encryption, such as elliptic curve cryptography (ECC), are used to secure transactions in modern blockchain platforms[Chandel2019AMA]. These techniques ensure that transactions are tamper-proof and can be validated by all participants in the network.
@@ 4.1 Basic Definition and Properties of Hash Functions
Hash functions play a fundamental role in many cryptographic systems, particularly in blockchain technology. At their core, a hash function is a mathematical algorithm that takes an input (or ”message”) and returns a fixed-size string of bytes, typically in the form of a hash value[Smart2016HashFM]. This output is unique to the specific input, but the same input will always result in the same hash value. For example, in blockchain systems, hash functions are used to maintain data integrity, verify transactions, and protect the system from manipulation. There are several essential properties of hash functions, two of which are particularly important: one-wayness and collision resistance.
@@ 4.1.1 One-Wayness and Irreversibility
One of the most critical properties of a cryptographic hash function is one-wayness[Boldyreva2009FoundationsON]. This means that, given the hash value, it is computationally infeasible to reverse-engineer the input that produced it. Essentially, hash functions are designed to be irreversible. Even if someone has the hash value, they cannot retrieve the original input data easily.
@@
This irreversibility is what protects data in blockchain systems[Wang2024BlockchainSA]. For example, even if someone obtains the hash of a block or transaction, they cannot use it to figure out the details of the original transaction.
@@ 4.1.2 Collision Resistance and Security
Another critical property of cryptographic hash functions is collision resistance[Stevens2021ASO]. A collision occurs when two different inputs produce the same hash output. A secure hash function is designed to make it computationally infeasible to find any two distinct inputs that result in the same hash value.
@@ 4.2 Simple Hash Algorithms
While cryptographic hash functions are designed with complex mathematics to ensure security, there are simpler forms of hash functions that are used in various applications, especially when high security is not the main concern. These simpler algorithms are often used for tasks[Wieder2017HashingLB] such as quick data indexing, load balancing, or data distribution across different storage locations.
@@ 4.2.1 Modulo Operation and Hashing
One of the most straightforward ways to generate a hash value is by using the modulo operation[Girault1987HashFunctionsUM]. In this approach, the hash value is obtained by dividing the input value by a certain number (called the modulus) and taking the remainder.